OpenStudy (anonymous):

What is the dot product of a = -2i - 5j and b = i - 4j? A. 18 B. -18 C. 22 D. -22 Please Help!

OpenStudy (surry99):

The general method for the dot product of two vectors is: a = a1i + a2j + a3k b= b1i + b2j + b3k Then a.b = a1*b1 + a2*b2 + a3*b3 note that when you take the dot product of two vectors you get a scalar Give it a try!

OpenStudy (anonymous):

I will, thank you so much!

OpenStudy (surry99):

Go ahead and try and I can check it for you if you like